OpenStudy (anonymous):

What is the torque generated at the elbow by a 90-N force pulling on the forearm at an angle of 110deg from the horizon at a point 18cm from the elbows axis of rotation. The forearm is postioned horizontal to the ground

OpenStudy (unklerhaukus):

\[\vec\tau=\vec r\times \vec F\] \[\tau=rF\sin\theta\]

OpenStudy (anonymous):

why did you choose sin? i think i just can't picture the diagram properly

OpenStudy (unklerhaukus):